OpenStudy (anonymous):

If a rancher sold 1/3 of his cattle and then sold another 1/4 of the original amount, what part of his cattle was sold?

OpenStudy (nirmalnema):

total cattle sold = 1/4 + 1/3 = 7/12 let the total no of cattles be x so no of cattles left.= x- 7/12 = 12x-7 / 12

OpenStudy (anonymous):

Harriet had a 2 1/4-foot length of board. She will use 1 2/3 feet for a project today. How many feet of board will remain? (Hint: you can use 9/4 and 5/3 if you wish.

OpenStudy (nirmalnema):

total length used=9/4+5/3=47/12 total length be x so remaning part is x- 47/12 = 12x-47 / 12

OpenStudy (nirmalnema):

do you understand or not..??